# Quadratic-time computations for pseudo-Anosov mapping classes

### v1

@@ 1. Introduction
Let $S=S_{g,p}$ be the surface obtained from the connected sum of $g$ tori by deleting $p$ points. The Nielsen–Thurston classification theorem says that the elements of the mapping class group $\operatorname{Mod}(S)$ fall into three categories: periodic, reducible, and pseudo-Anosov, and that the last category is exclusive from the other two. To a pseudo-Anosov element $f$ there are a two pieces of data associated: a pair of measured foliations on $S$ and a real number $\lambda>1$. These are called the stable and unstable foliations of $f$ and the stretch factor.
The main goal of this paper is to give a quadratic-time algorithm to compute the stretch factor and foliations for a pseudo-Anosov mapping class. The input for the algorithm is a product of generators of $\operatorname{Mod}(S)$ and the output is a pair $(\tau,D)$ where $\tau$ is a train track in $S$ and $D$ is a rational matrix acting on the vector space spanned by the branches of $\tau$. The stretch factor $\lambda$ is the largest real eigenvalue of $D$ and unstable foliation is given by any corresponding positive eigenvector. We refer to the pair $(\tau,D)$ as a pseudo-Anosov package for $f$.
@@
If the word length of the input is $n$, we show that the pseudo-Anosov package can be computed using $O(n^{2})$ binary operations. The details of the algorithm are given below, and the quadratic upper bound on the complexity is stated in Corollary 1.2. There is no other algorithm whose running time is known to be faster than exponential. Our paper does not address the issue of computing the eigendata for $D$, since there are many well-known (and fast) algorithms for this.
@@
We now explain the broad idea behind the algorithm. The group $\operatorname{Mod}(S)$ acts in a piecewise-linear fashion on $\operatorname{MF}(S)$, the space of measured foliations on $S$. Our main technical theorem, Theorem 1.1, says that there is a number $Q=Q(S)$, a cell structure on $\operatorname{MF}(S)$, and a point $c$ in $\operatorname{MF}(S)$ so that for any pseudo-Anosov $f\in\operatorname{Mod}(S)$ the point $f^{Q}(c)$ lies in a region of $\operatorname{MF}(S)$ that has two properties: (1) it contains the unstable foliation for $f$ and (2) the action of $f$ on the region is linear. From these two properties we may readily extract the desired pseudo-Anosov package, as the unstable foliation is an eigenvector of the linear map from item (2). We can compute $f^{Q}(c)$ and the action of $f$ at $f^{Q}(c)$ in quadratic time; essentially this reduces to the problem of multiplying together $Qn$ matrices from a given finite list of matrices, where $n$ is the word length of $f$.
@@
Our algorithm has its origins in work of Thurston and Mosher (as exposited by Casson–Bleiler [10, p. 146]), who gave a method for explicitly computing the action of $\operatorname{Mod}(S)$ on $\operatorname{MF}(S)$. In his famous research announcement, Thurston suggested that under iteration by a pseudo-Anosov mapping class, points in $\operatorname{MF}(S)$ converge to the unstable foliation “rather quickly,” Our Theorem 1.1 is an affirmation of this.
@@
Other works have taken advantage of the fact that the sequence $f^{k}(c)$ eventually—and quickly—lands in the “right” linear region for $f$; see specifically the work of Menzel–Parker [20], Moussafir [22], Finn–Thiffeault [13], Hall (who wrote the computer program Dynn) [16], the fourth-named author with Hall [17], and Bell (who wrote the computer programs flipper and curver) [6]. Our Theorem 1.1 is the first result to quantify how quickly the sequence $f^{k}(c)$ lands in the right linear region.
There are also algorithms to compute stretch factors and foliations that do not use the action on $\operatorname{MF}(S)$ directly. Most notably, Bestvina–Handel gave an algorithm that uses train tracks and is based on the Stallings theory of folding [7]. The expectation is that this algorithm has exponential time complexity. Brinkmann implemented the algorithm in the computer program Xtrain; see [9].
@@ 1.1. The main theorem
In this paper, a integral cone complex is a sort of cell complex where a cell is a positive cone in some $\mathbb{R}^{d}$ defined by a finite set of integer linear equations and integer inequalities; we will refer to such a cone as an integral linear cone. Such a cell has a finite set of faces, each being a cell in the same sense. Cells are glued along their faces, the gluing maps being linear.
With this basic definition in hand, we have several further definitions to make:
@@
    integral cone complex maps and linear regions,
@@
    the projectivized complex, vertex rays, induced maps, and radius,
@@
(b) If $Y$ is an integral cone complex, there is a projectivized version of $Y$, denoted $\mathcal{P}(Y)$, which is a cell complex in the usual sense. For instance the projectivized complex for $\operatorname{MF}(S)$ is $\mathcal{P}(\operatorname{MF}(S))=\operatorname{PMF}(S)$ (Thurston described the $\operatorname{Mod}(S)$-action on $\operatorname{PMF}(S)$ as opposed to $\operatorname{MF}(S)$). If $f:Y\to Y$ is an integral cone complex map, then there is an induced cellular map $\mathcal{P}(Y)\to\mathcal{P}(Y)$. Also, a vertex ray for a integral cone in an integral cone complex $Y$ is a vertex of the corresponding polyhedron in $\mathcal{P}(Y)$.
We now aim to define the radius of $Y$, or $\mathcal{P}(Y)$, as follows. First, we define a graph $\mathcal{G}_{Y}$ whose vertices are the top-dimensional cells $\mathcal{P}(Y)$ and whose edges connect cells of $\mathcal{P}(Y)$ that have nonempty intersection. For a set of vertices $W$ of $\mathcal{G}_{Y}$, we define the radius of $Y$ at $W$ to be the supremal distance in $\mathcal{G}_{Y}$ from a vertex of $\mathcal{G}_{Y}$ to $W$. For a vertex $c$ of $\mathcal{P}(Y)$, we define the radius of $Y$ at $c$ to be the radius of $Y$ at $W_{c}$, the set of top-dimensional cells of $\mathcal{P}(Y)$ containing $c$.
As an example, if $Y$ is any triangulation of $S^{2}$ given by the equator and any number of longitudes, then the radius of $Y$ with respect to the vertex at the north pole is 1. As we will see, there is an analogous situation for $\operatorname{MF}(S)$.
(c) Suppose that $G$ is a group that acts on the integral cone complex $Y$ by integral cone complex maps, and say that $G$ is generated by $\Gamma=\{h_{\ell}\}$. We denote by $Y_{\Gamma}$ the collection of integral cone complex subdivisions $\{Y_{h_{\ell}}\}$ (so there is one subdivision for each generator). As above, we denote the cells of the $h_{\ell}$-subdivision of $V_{i}$ by $V_{i,k}^{h_{\ell}}$.
@@
(d) Let $Y$ be an integral cone complex and let $f:Y\to Y$ be an integral cone complex map. For a point $v$ in $Y$ and a $\lambda\geq 0$, the point $\lambda v$ is a well-defined point of $Y$. With this in mind, a PL-eigenvector for $f$ is a point $v\in Y$ so that $f(v)=\lambda v$ for some $\lambda\geq 0$. We emphasize that if $A$ represents $f$ on a cone $V$ of $Y$, a real eigenvector $v$ for $A$ is not necessarily a PL-eigenvector for $f$. This is because $v$ might not lie in the cone of $\mathbb{R}^{d}$ corresponding to $V$.
(e) Next, a PL-eigenregion for the integral cone complex map $f$ is a linear piece $U$ for $f$ that contains a PL-eigenvector $v$. The linear transformation associated to $f|U$ will recognize $v$ as an eigenvector. More specifically, if $f|U$ has the triple $(A,V_{i},V_{j})$ with $i=j$ then $v$ will be a (classical) eigenvector for $A$. When $i\neq j$, we can choose coordinates on $V_{i}$ and $V_{j}$ that agree on the cell $V_{ij}=V_{i}\cap V_{j}$. In these coordinates, we again have that $v$ is an eigenvector for $A$.
@@
(f) Next, an integral cone complex map $f:Y\to Y$ is projectively source-sink if the induced cellular map $\mathcal{P}(Y)\to\mathcal{P}(Y)$ has source-sink dynamics. This means that the induced map has exactly two fixed points, the source and the sink, and that the $f$-iterates of all points outside of the source converge to the sink. In particular, the only two PL-eigenvectors for $f$ (up to scale) are representatives of the source and the sink.
(g) Let $Y$ be an integral cone complex and $f:Y\to Y$ a projectively source-sink integral cone complex map. Let $U$ be a PL-eigenregion for $f$ with triple $(A,V_{i},V_{j})$. There exists a maximal subcone $W$ of $V_{i}\cap V_{j}$ invariant under $f$, and the map $f|W$ is a dynamical map in the sense that $f$ maps $W$ to itself. The cone $W$ is nontrivial since it contains the ray corresponding to the sink. In Section 1.2 we give an algorithmic procedure to find $W$ and to compute the linear map associated to $f|W$. We also show that, in any $\mathbb{R}^{d}$-coordinates on $V_{i}$, the cone $W$ is the intersection of $V_{i}$ with a subspace of $\mathbb{R}^{d}$. Hence it makes sense to project $V_{i}$ onto $W$.
@@
There is a particularly convenient class of train track cell structures on $\operatorname{MF}(S)$, namely, the ones arising from Dehn–Thurston train tracks. These are defined as follows. First, we choose a pants decomposition of $S$; we will refer to the curves of the pants decomposition as base curves. From there, we decompose $S$ into a collection of annuli and pairs of pants. The core curves of the annuli are exactly the base curves and the pairs of pants are the closures of the complementary regions. Each pair of pants has between one and three boundary components, with the other holes accounted for by punctures. As in the book by Penner–Harer [24, §2.6], there are four model (sub)-train tracks on each annulus and at most four on each pair of pants. By gluing together all possible models, we obtain the collection of Dehn–Thurston train tracks. We prove in Lemma 7.2 that if $Y$ is a Dehn–Thurston cell structure on $\operatorname{MF}(S)$ and $c$ is one of the base curves, then the radius of $Y$ at $c$ is 1.
@@
Statement of the main theorem
Let $S=S_{g,p}$. For the statement of the main theorem, we must define a constant $Q=Q(S)$. We build up the definition of $Q$ by defining a sequence of preliminary constants. We do this in order to give a sense of how $Q$ arises, and to make feasible to improve our $Q$ by by improving the intermediate constants (and hence improve the implementation). In the end, we give a definition of $Q$ using the intermediary constants, and also a formula given directly in terms of $g$ and $p$; the impatient reader is encouraged to skip to the latter definition of $Q$.
@@
The number $K=K(S)$ is the maximum number of separatrices for a measured foliation on $S$; we refer to $K$ as the separatrix complexity of $S$.
@@
We may now state our main theorem in terms of Dehn–Thurston cell structures, Nielsen–Thurston eigenregions, and the number $Q=Q(S)$ just defined.

### v2

@@ 1. Introduction
In his famous survey paper from 1982, “Three dimensional manifolds, Kleinian groups and hyperbolic geometry,” William Thurston laid out a series of 24 problems that would shape the field of low-dimensional topology for the ensuing decades [30]. Problem 20 is:
Develop a computer program to calculate the canonical form for a general diffeomorphism of a surface, and to calculate the action of the group of diffeomorphisms on $P\!\mathscr{L}_{0}(S)$. Use this to implement an algorithm for the word problem and the conjugacy problem in the group of isotopy classes of diffeomorphisms of a surface.
Here, $P\!\mathscr{L}_{0}(S)$ is the space of (compactly supported) projective measured laminations on a surface $S$. One reason for Thurston’s interest in this question is revealed by his Problem 21: Develop a computer program to calculate hyperbolic structures on 3-manifolds. Because of the central place of the theory of surfaces in low-dimensional topology, Problem 20 has importance in its own right.
Significant progress has been made on Thurston’s problem (see below), but it has not been completely solved; there are algorithms to compute the canonical form in general, but not computer programs that work for all compact surfaces. In this paper we give, for any compact surface, an easily-implementable and efficient algorithm for computing the canonical form in the case where the diffeomorphism is assumed to be of a certain type (namely, pseudo-Anosov).
@@
To state our main result requires some setup. Let $S=S_{g,p}$ be the surface obtained from the connected sum of $g$ tori by deleting $p$ points. The Nielsen–Thurston classification theorem says that the elements of the mapping class group $\operatorname{Mod}(S)$ fall into three categories: periodic, reducible, and pseudo-Anosov, and that the last category is exclusive from the other two. To a pseudo-Anosov element $f$ there are a two pieces of data associated: a pair of measured foliations on $S$ and a real number $\lambda>1$. These are called the stable and unstable foliations of $f$ and the stretch factor. Any foliation in $S$ can be represented by a train track $\tau$ in $S$ along with a vector in the vector space $V(\tau)$ spanned by the branches of $\tau$.
A Nielsen–Thurston package for $f\in\operatorname{Mod}(S)$ is a pair $(\tau,D)$, where $\tau$ is train track in $S$ and $D$ is a rational matrix that represents a linear map of $V(\tau)$, that has the stretch factor of $f$ as its largest real eigenvalue, and that has the unstable foliation of $f$ as a corresponding positive eigenvector.
Main Theorem.
Fix a surface $S$ and let $\Gamma$ be any finite generating set for $\operatorname{Mod}(S)$. There is an explicit algorithm whose input is any word $w$ in $\Gamma$ that represents a pseudo-Anosov element of $\operatorname{Mod}(S)$, whose output is a Nielsen–Thurston package $(\tau,D)$ for that element, and whose running time is quadratic in the length of $w$.
This main theorem is stated more explicitly as Corollary 1.2 below. We consider this result to be a step towards solving Thurston’s problem in general. Our algorithm is in the spirit of Thurston’s problem: we directly compute (part of) the action of a mapping class on the space of measured foliations on a surface $S$, which is canonically isomorphic to $P\!\mathscr{L}_{0}(S)$.
@@
There is no other algorithm for computing stretch factors and foliations whose running time is known to be faster than exponential. Our paper does not address the issue of computing the eigendata for the matrix $D$, since there are many well-known (and fast) algorithms for this.
@@
We now explain the broad idea behind the algorithm. The group $\operatorname{Mod}(S)$ acts in a piecewise-linear fashion on $\operatorname{MF}(S)$, the space of measured foliations on $S$. Our main technical theorem, Theorem 1.1, says that there is a number $Q=Q(S)$, a cell structure on $\operatorname{MF}(S)$, and a point $c$ in $\operatorname{MF}(S)$ so that for any pseudo-Anosov $f\in\operatorname{Mod}(S)$ the point $f^{Q}(c)$ lies in a region of $\operatorname{MF}(S)$ that has two properties: (1) it contains the unstable foliation for $f$ and (2) the action of $f$ on the region is linear. From these two properties we may readily extract the desired Nielsen–Thurston package, as the unstable foliation is an eigenvector of the linear map from item (2). We can compute $f^{Q}(c)$ and the action of $f$ at $f^{Q}(c)$ in quadratic time; essentially this reduces to the problem of multiplying together $Qn$ matrices from a given finite list of matrices, where $n$ is the word length of $f$.
@@
Our algorithm has its origins in work of Thurston and Mosher (as exposited by Casson–Bleiler [10, p. 146]), who gave a method for explicitly computing the action of $\operatorname{Mod}(S)$ on $\operatorname{MF}(S)$. In his famous research announcement, Thurston suggested that under iteration by a pseudo-Anosov mapping class, points in $\operatorname{MF}(S)$ converge to the unstable foliation “rather quickly.” Our Theorem 1.1 is an affirmation of this.
@@
Other works have taken advantage of the fact that the sequence $f^{k}(c)$ eventually—and quickly—lands in the “right” linear region for $f$; see the work of Menzel–Parker [20], Moussafir [22], Finn–Thiffeault [13], Hall (who wrote the computer program Dynn) [16], the fourth-named author with Hall [17], and Bell (who wrote the computer programs flipper and curver) [6]. Our Theorem 1.1 is the first result to quantify how quickly the sequence $f^{k}(c)$ lands in the right linear region.
There are also algorithms to compute stretch factors and foliations that do not use the action on $\operatorname{MF}(S)$ directly. Most notably, Bestvina–Handel gave an algorithm that uses train tracks and is based on the Stallings theory of folding [7]. The expectation is that this algorithm has exponential time complexity. Brinkmann [9] implemented the algorithm for once-punctured surfaces in the computer program Xtrain and Hall implemented it for punctured disks in the program Trains.
@@ 1.1. The main technical result
In this paper, an integral cone complex is a sort of cell complex where a cell is a positive cone in some $\mathbb{R}^{d}$ defined as the set of solutions to a finite set of integer linear equations and integer inequalities; we will refer to such a cone as an integral linear cone. Such a cell has a finite set of faces, each being a cell in the same sense. Cells are glued along their faces, the gluing maps being linear.
Figure 1. A geometric realization of the integral cone complex $Y_{\Delta}$
We now present a basic example of an integral cone complex $Y_{\Delta}$ that we will treat as a running example. Let $V_{1}$, $V_{2}$, and $V_{3}$ each be the first quadrant in $\mathbb{R}^{2}$. We glue the face of $V_{i}$ on the $x$-axis to the face of $V_{i+1}$ on the $y$-axis (subscripts considered modulo 3) by the map $(t,0)\mapsto(0,t)$. The complex $Y_{\Delta}$ is homeomorphic to $\mathbb{R}^{2}$.
The integral cone complex $Y_{\Delta}$ can be geometrically realized in $\mathbb{R}^{3}$ as the boundary of the first octant (by a geometric realization, we mean an embedding into $\mathbb{R}^{N}$ mapping cells to cells). The complex $Y_{\Delta}$ is also geometrically realized as the set of points in the first octant of $\mathbb{R}^{3}$ whose coordinates satisfy a degenerate triangle inequality, meaning that one coordinate is the sum of the other two. See Figure 1 for an illustration.
With the definition of an integral cone complex in hand, we have several further definitions to make:
@@
    integral linear maps, integral cone complex maps, and linear regions,
@@
    the projectivized complex, vertex rays, and induced maps,
@@
To illustrate, we return to our running example $Y_{\Delta}$. Let
$V_{1,1}$ =$=\{(x,y)\in V_{1}\mid y\leq x\},\text{and}$
$V_{1,2}$ =$=\{(x,y)\in V_{1}\mid x\leq y\}.$
There is an integral cone complex map $f_{1}$ given by the four 4-tuples
$\left(A_{1,1},V_{1},V_{1},V_{1,1}\right),\left(A_{1,2},V_{1},V_{2},V_{1,2}\right),\left(A_{2},V_{2},V_{3},V_{2}\right),\text{ and }\left(A_{3},V_{3},V_{3},V_{3}\right),$
where
$A_{1,1}=\left(\begin{matrix}[r]1&-1\\ 0&1\end{matrix}\right),\ \ A_{1,2}=\left(\begin{matrix}[r]1&0\\ -1&1\end{matrix}\right),\ \ A_{2}=\left(\begin{matrix}[r]1&1\\ 0&1\end{matrix}\right),\ \ A_{3}=\left(\begin{matrix}[r]1&0\\ 1&1\end{matrix}\right),$
There is another cone complex map $r$ given by permuting the $V_{1}$ cyclically, mapping each to the next by the identity matrix.
(b) If $Y$ is an integral cone complex, there is a projectivized version of $Y$, denoted $\mathcal{P}(Y)$, which is a cell complex in the usual sense. For instance the projectivized complex for $\operatorname{MF}(S)$ is $\mathcal{P}(\operatorname{MF}(S))=\operatorname{PMF}(S)$ (Thurston described the $\operatorname{Mod}(S)$-action on $\operatorname{PMF}(S)$ as opposed to $\operatorname{MF}(S)$). If $f:Y\to Y$ is an integral cone complex map, then there is an induced cellular map $\mathcal{P}(Y_{1})\to\mathcal{P}(Y_{1})$, where $Y_{1}$ is an integral cone complex subdivision of $Y$. We also define a vertex ray of an integral cone complex $Y$ to be the ray corresponding to a vertex of $\mathcal{P}(Y)$.
For our running example $Y_{\Delta}$, the cell complex $\mathcal{P}(Y_{\Delta})$ is a triangle. Under the geometric realization given by Figure 1, $\mathcal{P}(Y_{\Delta})$ is identified with the triangle shown in the picture. There are three vertex rays in $Y_{\Delta}$. The subdivision associated to the map $f_{1}$ is given by adding a vertex to $\mathcal{P}(Y_{\Delta})$ corresponding to the ray spanned by $(1,1)$ in $V_{1}$; this is the midpoint of the $V_{1}$-edge of $\mathcal{P}(Y_{\Delta})$.
(c) We now set out to define the radius of an integral cone complex $Y$—which we also refer to as the radius of $\mathcal{P}(Y)$. First, we define a graph $\mathcal{G}_{Y}$ whose vertices are the top-dimensional cells of $\mathcal{P}(Y)$ and whose edges connect cells of $\mathcal{P}(Y)$ with nonempty intersection. For a set of vertices $W$ of $\mathcal{G}_{Y}$, we define the radius of $Y$ at $W$ to be the supremal distance in $\mathcal{G}_{Y}$ from a vertex of $\mathcal{G}_{Y}$ to $W$. For a vertex $c$ of $\mathcal{P}(Y)$, we define the radius of $Y$ at $c$ to be the radius of $Y$ at $W_{c}$, the set of top-dimensional cells of $\mathcal{P}(Y)$ containing $c$.
For the running example $Y_{\Delta}$, the graph $\mathcal{G}_{Y}$ is a triangle (dual to $\mathcal{P}(Y_{\Delta})$), and so the radius with respect to any $W$ is at most 1. As another example, if $Y$ is any triangulation of $S^{2}$ given by the equator and any number of longitudes, then the radius of $Y$ with respect to the vertex at the north pole is 1. As we will see, this example mirrors the situation for $\operatorname{MF}(S)$.
(d) Suppose that $G$ is a group that acts on the integral cone complex $Y$ by integral cone complex maps, and say that $G$ is generated by $\Gamma=\{h_{\ell}\}$. We denote by $Y_{\Gamma}$ the collection of integral cone complex subdivisions $\{Y_{h_{\ell}}\}$ (so there is one subdivision for each generator). As above, we denote the cells of the $h_{\ell}$-subdivision of $V_{i}$ by $V_{i,k}^{h_{\ell}}$.
@@
(e) Let $Y$ be an integral cone complex and let $f:Y\to Y$ be an integral cone complex map. For a point $v$ in $Y$ and a $\lambda\geq 0$, the point $\lambda v$ is a well-defined point of $Y$. With this in mind, a PL-eigenvector for $f$ is a point $v\in Y$ so that $f(v)=\lambda v$ for some $\lambda\geq 0$. We emphasize that if $A$ represents $f$ on a cone $V$ of $Y$, a real eigenvector $v$ for $A$ is not necessarily a PL-eigenvector for $f$. This is because $v$ might not lie in the cone of $\mathbb{R}^{d}$ corresponding to $V$.
In our running example, the point $(1,0)\in V_{1}$ is a PL-eigenvector of the integral cone complex map $f_{1}$ described above, since it is an eigenvector of the matrix $A_{1,1}$ and since the domain of this linear piece is contained in the codomain. By the same token, $(0,1)\in V_{1}$ is not a PL-eigenvector of $f_{1}$: it is an eigenvector of $A_{1,2}$ but its image is $(0,1)\in V_{2}$, which is not identified with (any multiple of) $(0,1)\in V_{1}$.
(f) Next, a PL-eigenregion for the integral cone complex map $f$ is a linear piece $U$ for $f$ that contains a PL-eigenvector $v$. The linear transformation associated to $f|U$ will recognize $v$ as an eigenvector. More specifically, if $f|U$ has the triple $(A,V_{i},V_{j})$ with $i=j$ then $v$ will be a (classical) eigenvector for $A$. When $i\neq j$, we can choose coordinates on $V_{i}$ and $V_{j}$ that agree on the cell $V_{ij}=V_{i}\cap V_{j}$. In these coordinates, we again have that $v$ is an eigenvector for $A$.
@@
In the running example, the map $f_{1}$ two PL-eigenregions for the PL-eigenvector $(1,0)\in V_{1}$, namely, $V_{1,1}$ and $V_{3}$. We already discussed the former region. Similarly, $V_{3}$ is a PL-eigenregion for $f_{1}$ since $(0,1)$ is an eigenvector for $A_{3}$ and the third piece of $f_{1}$ has the same domain and codomain, which contains $(0,1)$.
(g) Next, an integral cone complex map $f:Y\to Y$ is projectively source-sink if the induced cellular map $\mathcal{P}(Y)\to\mathcal{P}(Y)$ has source-sink dynamics. This means that the induced map has exactly two fixed points, the source and the sink, and that the $f$-iterates of all points outside of the source converge to the sink. In particular, the only two PL-eigenvectors for $f$ (up to scale) are representatives of the source and the sink.
In the running example, if we take $f_{2}=r\circ f_{1}\circ r^{-1}$, then the integral cone complex map
$f_{2}^{-1}\circ f_{1}$
is projectively source-sink. The sink and source correspond to the vectors $(\varphi,1)\in V_{3}$ and $(1,\varphi)\in V_{1}$, where $\varphi$ is the golden mean. The corresponding triples are
$(D,V_{3},V_{3})\text{ and }(E,V_{1},V_{1})$
where
$D=\left(\begin{matrix}[r]2&1\\ 1&1\end{matrix}\right)\text{ and }E=\left(\begin{matrix}[r]2&-1\\ -1&1\end{matrix}\right).$
(h) Let $Y$ be an integral cone complex and $f:Y\to Y$ a projectively source-sink integral cone complex map. Let $U$ be a PL-eigenregion for the sink of $f$ with triple $(A,V_{i},V_{j})$. There exists a maximal subcone $W$ of $V_{i}\cap V_{j}$ invariant under $f$, and the map $f|W$ is a dynamical map in the sense that $f$ maps $W$ to itself. The cone $W$ is nontrivial since it contains the ray corresponding to the sink. In Section 1.2 we give an algorithmic procedure to find $W$ and to compute the linear map associated to $f|W$. We also show that, in any $\mathbb{R}^{d}$-coordinates on $V_{i}$, the cone $W$ is the intersection of $V_{i}$ with a subspace of $\mathbb{R}^{d}$. Hence it makes sense to project $V_{i}$ onto $W$.
@@
In the running example, we have $i=j$ (this is the generic scenario); specifically $W=V_{i}=V_{j}=V_{3}$. As such, the sink package for the map $f_{2}^{-1}\circ f_{1}$ is $(V_{3},D)$, where $D$ is the matrix given above.
@@
As an example, let $S$ be $S_{0,4}$, or equivalently the thrice-punctured plane. The train tracks $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ shown in Figure 2 give a cell structure on $\operatorname{MF}(S_{0,4})$. The weights $x_{i}$ and $y_{i}$ on $\tau_{i}$ given an identification of each cell $V(\tau_{i})$ with the first quadrant in $\mathbb{R}^{2}$. The resulting integral cone complex structure on $\operatorname{MF}(S_{0,4})$ is isomorphic to our running example $Y_{\Delta}$. The geometric realization given in Figure 1 has a geometric description, namely, it is given by the geometric intersection numbers of a weighted train track with three rays to infinity that emanate from the three marked points.
[…]
Figure 2. Train tracks giving a cell structure on $\operatorname{MF}(S_{0,4})$
There is a convenient class of train track cell structures on $\operatorname{MF}(S)$ that applies to an arbitrary $S$, namely, the ones arising from Dehn–Thurston train tracks. These are defined as follows. First, we choose a pants decomposition of $S$; we will refer to the curves of the pants decomposition as base curves. From there, we decompose $S$ into a collection of annuli and pairs of pants. The core curves of the annuli are the base curves and the pairs of pants are the closures of the complementary regions. Each pair of pants has between one and three boundary components, with the other holes accounted for by punctures. As in the book by Penner–Harer [24, §2.6], there are four model (sub)-train tracks on each annulus and at most four on each pair of pants. By gluing together all possible models, we obtain the collection of Dehn–Thurston train tracks. We prove in Lemma 7.2 that if $Y$ is a Dehn–Thurston cell structure on $\operatorname{MF}(S)$ and $c$ is one of the base curves, then the radius of $Y$ at $c$ is 1.
@@
In our running example, the left-handed half-twist $H_{a}$ about the curve $a$ in Figure 2 (obtained by a clockwise $\pi$-rotation of the two leftmost marked points) is given by the map $f_{1}$ described above (using the above identification of $\operatorname{MF}(S_{0,4})$ with $Y_{\Delta}$), and similarly the half-twist $H_{b}$ is given by $f_{2}$. The mapping class $H_{b}^{-1}H_{a}$ is pseudo-Anosov with stretch factor $\varphi^{2}$. By the above calculation, the sink package for $H_{b}^{-1}H_{a}$ is $(D,V(\tau_{3}))$.
Statement of the main technical theorem
Let $S=S_{g,p}$. For the statement of Theorem 1.1, we must define a constant $Q=Q(S)$. We build up the definition of $Q$ by defining a sequence of preliminary constants. We do this in order to give a sense of how $Q$ arises, and to make feasible to improve our $Q$ by by improving the intermediate constants (and hence improve the implementation). In the end, we give a definition of $Q$ using the intermediary constants, and also a formula given directly in terms of $g$ and $p$; the impatient reader is encouraged to skip to the latter definition of $Q$.
@@
The number $K=K(S)$ is the maximum number of separatrices for a measured foliation on $S$; we refer to $K$ as the separatrix complexity of $S$. In other words, this is the maximum possible sum of the degrees of the singularities of a foliation on $S$.
@@
We may now state our main technical result in terms of Dehn–Thurston cell structures, Nielsen–Thurston eigenregions, and the number $Q=Q(S)$ just defined.